OpenStudy (anonymous):

Solve this using factoring a^2 + 8a - 20 = 0

OpenStudy (anonymous):

ConfusedStar... try www.wolframalpha.com enter what you just did into that.. easy solve..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(a-2) (a+10)

OpenStudy (anonymous):

what happen to the signs?

OpenStudy (anonymous):

a=10 a=-2

OpenStudy (anonymous):

opps thats a=-10 a=2

OpenStudy (anonymous):

a = -10 , 2

OpenStudy (anonymous):

(a-10) (a+2) ? this?

OpenStudy (anonymous):

(a+4)^2-36

OpenStudy (anonymous):

how does -10 became 10?

OpenStudy (anonymous):

and 2 became -2?

OpenStudy (anonymous):

a^2+8a-20=0 => (a-2)=0 1 8 -20 2 20 +________ 2 1 10 0 therefore, (a-2) & (a+10+0) => a=2 & a=-10

OpenStudy (anonymous):

The equation was factored into (a-2)(a+10)=0; This equation can be satisfied by either one of the terms of the LHS being equal to 0. So by equating each equal to 0; we get a=2,-10(a-2=0; a+10=0)